Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Conditional Probability

Expert replies
by savary » Mon Oct 12, 2015 1:52 am
I came across this question while studying the first chapter of the book (Introduction to Probability, Bertsekas 2nd 2008). The question is a solved example but I didn't understand it. I hope if someone can explain how to solve it rather than giving me the answer.

Q. A class consisting for 4 graduate and 12 undergraduate students is randomly divided into 4 groups of 4. What is the probability that each group includes a graduate student. [We interpret "randomly" to mean that given the assignment of some students to certain slots, any of the remaining students is equally likely to be assigned to any of the remaining slots]
Join the discussion
Source: — Problem Solving |

by GMATGuruNY » Mon Oct 12, 2015 2:59 am
savary wrote:A class consisting for 4 graduate and 12 undergraduate students is randomly divided into 4 groups of 4. What is the probability that each group includes a graduate student.
Let the 4 grad students be A, B, C and D.

Among the 4 groups of 4 are a total of 16 positions.
Student A can occupy any of the 16 positions.
.
Of the 15 remaining positions, 12 are not in A's group.
Thus:
P(B is not in A's group) = 12/15.

Of the 14 remaining positions, 8 are not in A's group or B's group.
Thus:
P(C is not in A's group or B's group) = 8/14.

Of the 13 remaining positions, 4 are not in A's group, B's group, or C's group.
Thus:
P(D is not in A's group, B's group, or C's group) = 4/13.

Since we want all of these events to happen, we MULTIPLY the probabilities:
12/15 * 8/14 * 4/13 = 64/455.
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion

by savary » Mon Oct 12, 2015 3:31 am
Thank you Mitch, this was very helpful. :)
Join the discussion

by Matt@VeritasPrep » Wed Oct 14, 2015 11:59 pm
I have this book on my shelf! (I can see why the worked example didn't help; the problem sets are great, but Bertsekas himself can be terse and tough to follow.)

Let me add a combinatorial approach, since this generalizes well to other problems you'll encounter later in real probability.

When we go to pick the first group, we want ONE grad and THREE undergrads. This can happen in 4*(12 choose 3) ways, since we're choosing one of four grads and three of twelve undergrads. This gives Target / Total = 4*(12 choose 3) / (16 choose 4). (The total is just all the ways we could randomly pick four people of the sixteen.)

When we go to pick the second group, we have 12 students left, and again want ONE grad and THREE undergrads. This can happen in 3*(9 choose 3) ways. This gives Target / Total = 3*(9 choose 3) / (12 choose 4)

When we go to pick the third group, similar logic gives us 2*(6 choose 3) ways. This gives Target / Total = 2*(6 choose 3) / (8 choose 4).

Since the fourth group is now predetermined, we don't have to bother with it.

Multiplying all the probabilities above gives us

4 * 3 * 2 * (12! / 3!9!) * (9! / 3!6!) * (6! / 3!3!) / ((16! / 4!12!) * (12! / 4!8!) * (8! / 4!4!))

As you can see when you write it out, this cleans up much more nicely than you'd expect, and leaves 64/455.
Join the discussion

by savary » Thu Dec 03, 2015 4:51 am
Matt, I'm sorry I didn't read your post earlier. I was going through the problems again after studying combinatorics. Actually, I was using the book along with video lectures from MIT OCW course entitled "Probability Models and Axioms". Having done the problem both ways, I found the combinatorial approach easier to follow but more time-consuming in calculations.
Join the discussion